# Fairness-Aware Multi-view Evidential Learning with Adaptive Prior

### v1

@@ Abstract
Multi-view evidential learning aims to integrate information from multiple views to improve prediction performance and provide trustworthy uncertainty esitimation. Most previous methods assume that view-specific evidence learning is naturally reliable. However, in practice, the evidence learning process tends to be biased. Through empirical analysis on real-world data, we reveal that samples tend to be assigned more evidence to support data-rich classes, thereby leading to unreliable uncertainty estimation in predictions. This motivates us to delve into a new Biased Evidential Multi-view Learning (BEML) problem. To this end, we propose Fairness-Aware Multi-view Evidential Learning (FAML). FAML first introduces an adaptive prior based on training trajectory, which acts as a regularization strategy to flexibly calibrate the biased evidence learning process. Furthermore, we explicitly incorporate a fairness constraint based on class-wise evidence variance to promote balanced evidence allocation. In the multi-view fusion stage, we propose an opinion alignment mechanism to mitigate view-specific bias across views, thereby encouraging the integration of consistent and mutually supportive evidence. Extensive experiments on five real-world multi-view datasets demonstrate that FAML achieves more balanced evidence allocation and improves both prediction performance and the reliability of uncertainty estimation compared to state-of-the-art methods.
@@ Introduction
Definition 1 Fairness Degree. Given the set of evidence vectors produced by Evidential Deep Neural Networks across all samples, the fairness degree is defined as the variance of the class-wise average evidence:
@@
Definition 2 Weighted Evidence Aggregation. Given two opinions $\mathcal{O}^{A}$ and $\mathcal{O}^{B}$ over an instance, let $e^{A}_{k}$, $e^{B}_{k}$ and $e_{k}$ represent the evidence for the $k$-th category from view $A$, view $B$, and the aggregated view, we have:
@@
The model optimization is elaborated in Algorithm 1 (Technical Appendix).
Theoretical Studies and Discussion
In this section, we provide theoretical insights to further justify the effectiveness of our adaptive training-trajectory-based prior in mitigating quantity-induced unfairness in multi-view evidential learning. The following propositions summarize the key theoretical characteristics of our approach. Proofs are provided in the Technical Appendix.
Proposition 1. The adaptive training-trajectory-based prior is a generalized form of the uniform prior employed by standard EDL. In ideal training conditions, it progressively converges to the uniform prior.
Based on Proposition 1, our adaptive prior adaptively adjust according to the predictive performance throughout the training process. This proposition highlights the theoretical consistency of our method, ensuring that the adaptive mechanism to be grounded theoretically in Subjective Logic (SL).
Proposition 2. The adaptive prior implicitly reformulates the base rate in the projected probability of a subjective opinion, leading to a biased uncertainty allocation that increases the predictive probability mass to underperforming classes.
Existing Multi-view Evidential Learning methods focus on designing fusion strategies to combine predictions from multiple views, they often overlook whether the evidence collection process is fairness or not. In practice, evidence learning often exhibits quantity-induced unfairness, where samples predicted as data-rich classes tend to be associated with low uncertainty, regardless of whether the prediction is actually correct. When incorporating a high-confidence but incorrect prediction from views, the fusion tends to favor this misleading opinion, overpowering low-confidence yet correct predictions from other views. As a result, the model is prone to producing confident yet incorrect decision.
@@
Ours 87.1 ± 0.9 99.5 ± 0.9 90.4 ± 1.5 72.8 ± 2.6 16.1 ± 3.7 23.8 ± 2.2 22.8 ± 5.4 27.8 ± 3.4
Table 1: Performance comparison in terms of ACC(%) and ECE(%) on test sets, with our method achieving state-of-the art performance. ± indicates the standard deviation for 5 random seeds, with the best mean scores highlighted in bold .
@@
To validate the effectiveness of FAML, we conduct experiments on five widely used real-word multi-view datasets as follows: Handwritten [^1] https://archive.ics.uci.edu/dataset/72/multiple+features is a six-view dataset, which comprises 2,000 instances of handwritten numerals ranging from ’0’ to ’9’. Animal222 https://www.vision.caltech.edu/Image contains 2 views and 10,158 images from 50 categories, utilizing two types of deep features extracted using DEACA and VGG. Scene15333https://doi.org/10.6084/m9.figshare.7007177.v1 includes 4485 images from 15 indoor and outdoor scene categories. We extract three types of features GIST, PHOG, and LBP. YaleB [^4] https://vision.ucsd.edu/content/yale-face-database consists of 650 facial images from 10 categories with three views Caltech101555https://data.caltech.edu/records/mzrjq-6wc02 comprises 8677 images from 101 classes. We select 2386 samples from 20 classes. We extract Gabor, Wavelet Moments, CENTRIST, HOG, GIST, and LBP as 6 views.

### v2

@@ Abstract
Multi-view evidential learning aims to integrate information from multiple views to improve prediction performance and provide trustworthy uncertainty estimation. Most previous methods assume that view-specific evidence learning is naturally reliable. However, in practice, the evidence learning process tends to be biased. Through empirical analysis on real-world data, we reveal that samples tend to be assigned more evidence to support data-rich classes, thereby leading to unreliable uncertainty estimation in predictions. This motivates us to delve into a new Biased Evidential Multi-view Learning (BEML) problem. To this end, we propose Fairness-Aware Multi-view Evidential Learning (FAML). FAML first introduces an adaptive prior based on training trajectories, which acts as a regularization strategy to flexibly calibrate the biased evidence learning process. Furthermore, we explicitly incorporate a fairness constraint based on class-wise evidence variance to promote balanced evidence allocation. In the multi-view fusion stage, we propose an opinion alignment mechanism to mitigate view-specific bias across views, thereby encouraging the integration of consistent and mutually supportive evidence. Theoretical analysis shows that FAML enhances fairness in the evidence learning process. Extensive experiments on six real-world multi-view datasets demonstrate that FAML achieves more balanced evidence allocation and improves both prediction performance and the reliability of uncertainty estimation compared to state-of-the-art methods.
@@ Introduction
Definition: Fairness Degree. Given the set of evidence vectors produced by Evidential Deep Neural Networks across all samples, the fairness degree is defined as the variance of the class-wise average evidence:
@@
Definition: Weighted Evidence Aggregation. Given two opinions $\mathcal{O}^{A}$ and $\mathcal{O}^{B}$ over an instance, let $e^{A}_{k}$, $e^{B}_{k}$ and $e_{k}$ represent the evidence for the $k$-th category from view $A$, view $B$, and the aggregated view, we have:
@@
Theoretical Studies
In this section, we aim to establish the generalization guarantees of the FAML framework, particularly for minority classes. The analysis is grounded in the margin theory of statistical learning, which connects the margin distribution achieved on the training set to the model’s generalization error. We first introduce the definition of Evidence Margin.
Definition 1 (Evidence Margin).
Let $h_{\theta}:\mathcal{X}\to\mathbb{R}^{K}_{+}$ denote a model, parameterized by $\theta$, that maps a sample $x_{n}$ to a $K$-dimensional non-negative evidence vector $e_{n}=[e_{n1},\ldots,e_{nK}]$. The evidence margin $\rho_{n}(\theta)$ of $x_{n}$ is then defined as
$\rho_{n}(\theta)=e_{nk}-\max_{j\neq k}\{e_{nj}\},$ (15)
where $k$ is the ground-truth class label of $x_{n}$.
To formally establish the connection between such a sample-wise margin and the model’s overall generalization ability, we now introduce a foundational result from statistical learning theory. The following lemma provides a general bound that relates the margin distribution achieved on the training set to the upper bound of the generalization error.
Lemma 2 (Margin Generalization Bound).
For a hypothesis space $\mathcal{H}$, given any margin $\rho>0$, with probability at least $1-\delta$, for all $h\in\mathcal{H}$, the true risk $R(h)$ satisfies:
$R(h)\leq\hat{R}_{S,\rho}(h)+C_{1}\sqrt{\frac{\text{Complexity}(\mathcal{H})}{N\rho^{2}}}+C_{2}\sqrt{\frac{\ln(1/\delta)}{N}}$ (16)
where $\hat{R}_{S,\rho}(h)$ is the proportion of samples in the training set $S$ with a margin less than $\rho$, $\text{Complexity}(\mathcal{H})$ is a complexity measure of the hypothesis space, and $C_{1},C_{2}$ are constants.
Lemma 2 shows that increasing the margin $\rho$ tightens the generalization bound via the $O(1/\rho)$ term. This allows us to use the change in $\rho$ as a bridge: by analyzing the effect of the FAML algorithm on $\rho$, we can explicitly determine its impact on the generalization error bound.
Theorem 3.
Let $h_{\theta}:\mathcal{X}\rightarrow\mathbb{R}_{+}^{K}$ be a model trained on a dataset $S$ of size $N$. For a minority class $k$ with imbalance ratio $\xi_{k}=N_{-k}/N_{k}\gg 1$, where $N_{k}$ and $N_{-k}$ are the number of samples belonging to class $k$ and other classes respectively, and under the condition that the adaptive prior $\beta_{k}$ increases when class $k$ performs poorly, minimizing the FAML objective $L(\theta)$ yields the following theoretical guarantees for the evidence margin $\rho_{n}(\theta)=e_{nk}-\max_{j\neq k}e_{nj}$:
  • Margin Increase for Minority Class. The expected margin improvement for class $k$ is bounded below by:
    $\mathbb{E}[\Delta\rho_{n}\mid y_{n}=k]\gtrsim\eta\psi^{\prime}\left(S_{\hat{\alpha}}\right)\cdot\left(\xi_{k}\cdot\frac{\Delta\beta_{k}}{S_{\hat{\alpha}}}-O\left(\frac{1}{\hat{\alpha}_{nk}^{2}}\right)\right),$
    where $\eta$ is the learning rate, $S_{\hat{\alpha}}=\sum_{j}\hat{\alpha}_{nj}$, $\hat{\alpha}_{nj}=e_{nj}+\beta_{j}$, and $\Delta\beta_{k}>0$ is the increase in the adaptive prior. For $\xi_{k}\gg 1$, the positive term dominates, ensuring a net margin increase.
  • Tighter Generalization Bound. The increase in margin tightens the generalization bound. The effective margin $\rho_{\textup{eff}}$ satisfies:
    $\rho_{\textup{eff}}\gtrsim\rho_{0}\left(1+\lambda\cdot\xi_{k}\cdot\Delta\beta_{k}\right),$
    where $\rho_{0}$ is the baseline margin and $\lambda>0$ is a scaling factor. This leads to an improved generalization error bound for the minority class by a factor of $\tilde{O}\left(1/\sqrt{\xi_{k}\Delta\beta_{k}}\right)$.
Theorem 3 delineates how the adaptive prior enhances the evidence margin and tightens the generalization bound, with particular efficacy for minority classes. First, as captured by the lower bound on the expected margin improvement, the corrective mechanism scales with the class imbalance ratio $\xi_{k}$. This scaling property ensures that the model automatically allocates greater corrective effort to classes that are both under-represented and under-performing. Second, the generalization improvement factor of $\tilde{O}(1/\sqrt{\xi_{k}\Delta\beta_{k}})$ quantifies this behavior: it shows that greater data imbalance (larger $\xi_{k}$) coupled with stronger model self-feedback (larger $\Delta\beta_{k}$) leads to greater generalization gain. This result provides a theoretical foundation for the superior performance of our method on long-tailed distributions.
@@
Table 1: Performance comparison in terms of ACC(%) and ECE(%) on test sets, ± indicates the standard deviation for 5 random seeds, with the best mean scores highlighted in bold.
@@
Ours 83.6 ± 1.0 99.5 ± 0.9 88.1 ± 1.5 67.8 ± 2.8 14.1 ± 1.3 11.4 ± 4.2 17.2 ± 1.5 26.9 ± 3.4
@@
To validate the effectiveness of FAML, we conduct experiments on five widely used real-word multi-view datasets as follows: Handwritten [^1] https://archive.ics.uci.edu/dataset/72/multiple+features is a six-view dataset, which comprises 2,000 instances of handwritten numerals ranging from ’0’ to ’9’. Animal [^2] https://www.vision.caltech.edu/Image contains 2 views and 10,158 images from 50 categories, utilizing two types of deep features extracted using DEACA and VGG. Scene15333https://doi.org/10.6084/m9.figshare.7007177.v1 includes 4485 images from 15 indoor and outdoor scene categories. We extract three types of features GIST, PHOG, and LBP. YaleB [^4] https://vision.ucsd.edu/content/yale-face-database consists of 650 facial images from 10 categories with three views Caltech101555https://data.caltech.edu/records/mzrjq-6wc02 comprises 8677 images from 101 classes. We select 2386 samples from 20 classes. We extract Gabor, Wavelet Moments, CENTRIST, HOG, GIST, and LBP as 6 views.